Based on previous research, an academic advisor predicted that tardiness would decrease if students were asked...

  1. Based on previous research, an academic advisor predicted that tardiness would decrease if students were asked to text their academic counselor as soon as they got to class. A sample of 25 students was selected for the study, and their attendance was monitored for 6 weeks. For the first 3 weeks, the students were asked to text as soon as they got to class. The next 3 weeks, students no longer needed to text. The table shows how many minutes students were late to class in the text and no text conditions, as well as the difference between these conditions for each student.
    1. What type of test is appropriate to analyze this data? (1pt)
    1. State the null and alternative hypotheses for a one-tailed test. (3pt)
    1. Calculate the appropriate test statisti (3pt)
    1. Report df and the critical value for a one-tailed test with alpha = .05. (2pt)
    1. Calculate Cohen’s d. (3pt)
    1. What percentage of the variance in tardiness is accounted for by the effect of texting? (3pt)
    1. Report the results (including d and r2) in APA format. (5pt)

      text

      no text

      difference

      12

      27

      15

      4

      24

      20

      7

      23

      16

      2

      5

      3

      2

      0

      -2

      3

      0

      -3

      2

      20

      18

      6

      2

      -4

      12

      20

      8

      8

      11

      3

      5

      6

      1

      18

      10

      -8

      24

      13

      -11

      20

      16

      -4

      21

      18

      -3

      2

      30

      28

      9

      2

      -7

      11

      15

      4

      15

      21

      6

      3

      23

      20

      14

      3

      -11

      12

      12

      0

      12

      10

      -2

      16

      19

      3

      10

      20

      10

      M = 10

      M = 14

      M = 4

      s = 6.519

      s = 8.855

      s = 10.5
Answer #1

1)

Paired t test

2)

Ho :   µd=   0                  
Ha :   µd <   0                  
                          

3)

  
                          
sample size ,    n =    25                  
                          
mean of sample 1,    x̅1=   10.000                  
                          
mean of sample 2,    x̅2=   14.000                  
                          
mean of difference ,    D̅ =ΣDi / n =   -4.0000                  
                          
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    10.5000                  
                          
std error , SE = Sd / √n =    10.5000   / √   25   =   2.1000      
                          
t-statistic = (D̅ - µd)/SE = (   -4   -   0   ) /    2.1000   =   -1.905

4)
                          
Degree of freedom, DF=   n - 1 =    24                  
t-critical value , t* =        --1.7109 [excel function: =t.inv(α,df) ]              
                          

5)


cohen's d = |Dbar/ std dev| =    0.381

                   

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!